OpenStudy (anonymous):

hich is less than zero point five five? seven-tenths three-fifths nine-twentieths sixteen twenty-fifths

OpenStudy (nurali):

Seven-tenths = 0.7 three-fifths = 3/5 = 0.6 nine-twentieths = 9/20 = 0.45 sixteen twenty-fifths = 16/25 = 0.64 The only one that is less than 0.55 is the second one, nine-twentieths = 9/20 = 0.45